OpenStudy (anonymous):

What mass, in grams, of fluorine gas (F2) is contained in a 10.3 liter tank at 29.8 degrees Celsius and 1.23 atmospheres? HELP PLEASE

OpenStudy (anonymous):

You will be using: PV = nRT, the Ideal Gas Law, with a few unit conversions. Does this sound familiar?

OpenStudy (anonymous):

Looking for mass in grams, we'll first solve for number of moles and then convert. Also, convert temperature units: 29.8 Celsius + 273.15 = 302.95 Kelvin PV = nRT, where P (Pressure) = 1.23 atmospheres where V (Volume) = 10.3 liters where n (moles) = ? where R (Universal Gas Constant) = 0.08206 Latm/molK where T (Temperature) = 29.8 Celsius + 273.15 = 302.95 Kelvin (we need to convert into Kelvin otherwise the formula doesn't work) n = PV/RT n = (1.23 atm)(10.3 L) / (0.08206 Latm/molK)(302.95 K) n = 0.50961 moles m = (n)(mu) where m = mass (in grams) where n = number of moles where mu = molar mass of F2 = 19 + 19 = 38 grams/mole mass = (0.50961 moles)(38 grams/mole) mass = 19.36527 ~ 19.3 grams